### v1

# Systematic Classification of History-Dependent Invariants in the Lorenz System: Regularization Classes and Dynamical Signatures

Abstract
Building upon the recent discovery of history-dependent dynamical invariants in the Lorenz system, this work presents a systematic exploration of a complete family of such conserved quantities. Through analysis of all 24 permutations in the augmented phase space, we identify eighteen distinct invariant structures ($K_{1}$$K_{18}$) and demonstrate that six permutations yield null results. A master table associates each invariant with its generating permutation, revealing that the class structure emerges naturally from the permutation ordering: Class I ($K_{1}$$K_{6}$, permutations 1abc), Class II ($K_{7}$$K_{12}$, permutations 2abc), and Class III ($K_{13}$$K_{18}$, permutations 3abc). The null permutations 4abc correspond to placing the auxiliary variable first, which yields no valid invariants due to algebraic consistency requirements imposed by Schwarz integrability conditions. A complete regularization procedure is developed for all three classes, with each class characterized by a distinct polynomial factor that removes singularities. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$; while the original variable satisfies a linear ODE with $u_{n}$-dependent coefficients, the regularized variable $v_{n}=p\cdot u_{n}$ evolves according to $\dot{v}_{n}=Q_{n}(x,y,z)$, independent of $v_{n}$ itself. Different conservation laws thereby impose distinct measures of trajectory history. Analysis of the functional independence structure establishes that any three invariants selected from different regularization classes provide independent constraints on trajectory histories, with representative invariants from each class defining a canonical “Triad” that characterizes the orbit structure. Statistical analysis of the auxiliary evolution functions $Q_{n}(t)$ reveals that invariants from different regularization classes exhibit distinct dynamical signatures: Class III invariants, whose polynomial part $P_{\mathrm{III}}$ contains terms involving the bilinear product $xy$, display significantly higher intermittency (kurtosis $\kappa_{\mathrm{III}}\approx 15$ versus $\kappa_{\mathrm{I}}\approx 8$ for Class I) and stronger asymmetry (skewness $|S_{\mathrm{III}}|\approx 2.7$ versus $|S_{\mathrm{I}}|\approx 1.6$), correlating with lobe-switching events where the product $xy$ undergoes rapid sign reversal. This geometric sensitivity establishes Class III invariants as probes of topological transitions. The class-dependent divergence of auxiliary evolution functions provides a quantitative measure of how different conservation principles impose incompatible demands on trajectory history, with the variation required to maintain each conservation law encoded in the statistics of the corresponding $\dot{v}_{n}$. High-precision numerical validation confirms conservation of all identified quantities.
I Introduction
The recent demonstration that the Lorenz system possesses history-dependent dynamical invariants has expanded the understanding of conservation laws in dissipative chaotic systems [1]. That work established the existence of a non-local conserved quantity whose constancy along trajectories arises through accumulation of the system’s past via integral terms involving an auxiliary variable. The construction revealed structural organization within an augmented phase space and provided a characterization of unstable periodic orbits as level sets of this conserved functional.
The construction of auxiliary variables to achieve conservation in non-conservative systems connects to several theoretical frameworks. In statistical mechanics, the Mori-Zwanzig formalism demonstrates that projecting high-dimensional dynamics onto lower-dimensional subspaces generates memory kernels, which are convolution integrals over the system’s history [2, 3]. The auxiliary variable $u(t)$ introduced for the Lorenz system represents a deterministic realization of the memory term that emerges when closing an intrinsically dissipative system. This construction aligns with the principles of dimension reduction in dynamical systems.
Independently, Hojman [5] demonstrated that conservation laws can be constructed without recourse to Lagrangian or Hamiltonian formulations, establishing that the connection between symmetries and conserved quantities extends beyond the variational framework. The present construction shares this non-Noetherian character: the invariants $K_{n}$ do not arise from continuous symmetries of an action functional but from algebraic compatibility between the Lorenz flow and the orthogonality ansatz.
However, the constructive method admits multiple distinct choices in the orthogonality ansatz used to augment the three-dimensional Lorenz phase space. The original work noted that “the constructive method employed can yield other conserved quantities,” deferring systematic exploration. This raises specific questions: How many independent history-dependent invariants exist? What determines their mathematical form? Do they possess distinct properties corresponding to different dynamical interpretations?
The present work addresses these questions through systematic analysis. By considering all 24 permutations of the phase space coordinates in the orthogonality construction, we identify exactly eighteen valid invariant structures and demonstrate that six permutations yield null results. The first coordinate in the permutation determines the regularization class. This provides both complete enumeration and structural understanding of the invariant family.
An important insight concerns the nature of the auxiliary variables. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$ satisfying a specific evolution equation $\dot{u}_{n}=f_{n}(x,y,z)$. A distinction must be drawn between two formulations: the original auxiliary variable $u_{n}(t)$, which diverges at nullcline crossings where the regularization polynomial vanishes, and the regularized variable $v_{n}(t)=p_{n}(x,y,z)\cdot u_{n}(t)$, which remains bounded for all time. The regularized formulation, in which each invariant takes the form $K_{n}=P_{n}(x,y,z)+v_{n}$, provides the mathematically proper framework. The fluctuation patterns of the regularized variables differ across classes: different invariants exhibit distinct dynamical signatures as they respond to the trajectory’s passage through various regions of phase space. Different conservation laws impose distinct measures of trajectory history. The kinematic variation required to maintain conservation of $K_{n}$ is encoded in the evolution function $Q_{n}(t)=\dot{v}_{n}$, and this variation differs across regularization classes.
The existence of the null class (Theorem 1) demonstrates that not all polynomial structures are compatible with the Lorenz flow: six of 24 permutations fail to yield valid invariants because Schwarz integrability conditions reduce to algebraic contradictions. This construction yields two principal results: the classification of compatible polynomial structures, which encodes information specific to the Lorenz dynamics, and the distinct dynamical signatures exhibited by different regularization classes, which provide complementary probes of attractor geometry.
The invariant $K$ presented in Ref. [1] was derived using an orthogonality ansatz of the form $\nabla C=(\epsilon_{1}\dot{\pi}_{1},\epsilon_{2}\dot{\pi}_{2},\epsilon_{3}\dot{\pi}_{3},\epsilon_{4}\dot{\pi}_{4})$, where each sign $\epsilon_{i}\in\{+1,-1\}$ is assigned independently to each permuted component. This general construction admits $24\times 2^{4}/2=192$ distinct ansätze, of which only 24 correspond to a uniform global sign. The present work focuses on the latter class, which yields the organizational structure described above. Both families share the same regularization polynomials, a connection suggesting underlying structural unity.
We also demonstrate that invariants from different regularization classes exhibit distinct dynamical signatures. Statistical analysis reveals different intermittency patterns in the chaotic regime: Class III invariants, whose polynomial parts $P_{\mathrm{III}}$ contain terms involving the bilinear product $xy$, exhibit enhanced sensitivity to lobe-switching events where this product changes sign. The rapid variation of $xy$ at the separatrix ($x=0$) forces correspondingly rapid variations in $Q_{\mathrm{III}}=-\dot{P}_{\mathrm{III}}$. These events represent topological transitions in the attractor’s symbolic dynamics [6]. The Class III evolution function $Q_{\mathrm{III}}(t)$ functions as a geometric probe that signals each addition to the trajectory’s symbolic sequence, establishing a quantitative link between continuous dynamics and discrete topological structure.
This paper is organized as follows. Section II reviews the constructive orthogonality method and introduces the permutation-based systematic exploration. Section III presents the complete family of eighteen invariants organized by permutation class, including regularization and the proof that permutations beginning with the auxiliary variable yield null results. Section IV establishes the functional independence structure and clarifies the role of invariant-specific auxiliary variables. Section V develops the geometric interpretation of history-dependent constraints. Section VI analyzes transformation properties under the discrete $\mathbb{Z}_{2}$ symmetry. Section VII provides numerical verification and statistical characterization. Section VIII synthesizes the results, including analysis of the physical interpretation of the invariants and their operational meaning.
II Systematic Extension of the Constructive Method
II.1 Review of the Orthogonality Approach
The Lorenz system is defined by the three coupled nonlinear ordinary differential equations
$\dot{x}$ $=\sigma(y-x),$ (1)
$\dot{y}$ $=x(\rho-z)-y,$ (2)
$\dot{z}$ $=xy-\beta z,$ (3)
where $\sigma$, $\rho$, and $\beta$ are positive parameters. For the classical parameter values $\sigma=10$, $\rho=28$, and $\beta=8/3$, the system exhibits chaotic dynamics on the Lorenz attractor [7, 8].
To construct conserved quantities for this odd-dimensional system, the phase space is augmented with an auxiliary variable $u$, yielding a four-dimensional state space $(x,y,z,u)$. A candidate conserved quantity $C(x,y,z,u)$ must satisfy
$\frac{dC}{dt}=\nabla C\cdot\mathbf{f}=0,$ (4)
where $\nabla\equiv(\partial_{x},\partial_{y},\partial_{z},\partial_{u})$, the flow vector is $\mathbf{f}\equiv(\dot{x},\dot{y},\dot{z},\dot{u})$, and $\dot{u}=f(x,y,z)$ is to be determined self-consistently.
II.2 Connection to the Mori-Zwanzig Formalism
The introduction of the auxiliary variable $u(t)$ admits a natural interpretation within the framework of non-equilibrium statistical mechanics. The Mori-Zwanzig projection operator formalism [2, 3] establishes that when a high-dimensional system is projected onto a lower-dimensional subspace, the resulting equations necessarily acquire memory terms:
$\frac{dA}{dt}=\Omega A+\int_{0}^{t}K(t-s)A(s)\,ds+F(t),$ (5)
where $K(t)$ is a memory kernel encoding the influence of eliminated degrees of freedom and $F(t)$ represents a fluctuating force orthogonal to the resolved variables.
The Lorenz system, derived from a low-order truncation of the Navier-Stokes equations, represents such a projection. The regularized auxiliary variable $v_{n}(t)$, which evolves according to
$v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}(x(s),y(s),z(s))\,ds,$ (6)
can be understood as a deterministic resolution of the memory kernel: it explicitly tracks the accumulated effect of the trajectory’s history that would otherwise manifest as non-Markovian dynamics. (The original auxiliary variable $u_{n}=v_{n}/p$ satisfies a more complex linear ODE with state-dependent coefficients; see Sec. III.2.)
A distinction must be drawn between two different approaches to treating unresolved degrees of freedom. In the Optimal Prediction framework of Chorin and collaborators [2], the fluctuating force $F(t)$ is treated stochastically: its effects are averaged over an ensemble, yielding a coarse-grained, probabilistic description. The fluctuation-dissipation theorem [3] then relates the memory kernel to the autocorrelation of this noise, establishing a thermodynamic connection between dissipation and fluctuations.
In contrast, the present construction constitutes a deterministic embedding: by extending the phase space with $v_{n}(t)$, we effectively absorb the orthogonal fluctuations into a resolved variable, achieving an exact closure in which $F(t)\equiv 0$ along any trajectory. This distinction has important consequences. The regularized variable $v_{n}(t)$ in this construction lacks the thermodynamic content of a true memory kernel in the Mori-Zwanzig sense: it does not satisfy a fluctuation-dissipation relation, and its “memory” is purely kinematic rather than thermodynamic. The variable functions as a deterministic history accumulator, recording the trajectory’s past without the statistical properties that would connect it to heat dissipation or entropy production.
This perspective clarifies what the construction achieves: it provides a geometric embedding that renders dissipative dynamics conservative in an extended space, at the cost of introducing a non-local (history-dependent) degree of freedom. The existence of exactly eighteen conserved quantities represents a structural property of the Lorenz equations that is not guaranteed a priori.
The paradigm for embedding dissipative systems within larger conservative frameworks has been developed systematically by Morrison [4], who demonstrated that many dissipative systems admit Hamiltonian formulations in extended phase spaces. The present construction represents a concrete realization of this principle for the Lorenz system.
Physical scaling. In the original Rayleigh-Bénard convection context from which the Lorenz equations derive [7], the state variables represent amplitudes of truncated Fourier modes: $x$ corresponds to the convective velocity mode, while $y$ and $z$ correspond to temperature perturbation modes. The parameters $\sigma$, $\rho$, and $\beta$ are dimensionless after appropriate scaling by the thermal diffusivity, cell height, and critical Rayleigh number. In this dimensionless formulation, the regularized auxiliary variable $v_{n}(t)$ inherits a kinematic interpretation as a generalized potential that accumulates historical effects of the modal amplitudes. For each regularization class, the polynomial factor $p_{n}(x,y,z)$ combines the physical modes in a specific manner: $p_{\mathrm{I}}=y-x$ couples velocity and temperature perturbations, $p_{\mathrm{II}}=y+x(z-\rho)$ involves the departure from criticality, and $p_{\mathrm{III}}=xy-\beta z$ couples to the nonlinear heat flux term. The invariant $K_{n}=P_{n}+v_{n}$ thereby encodes a balance between instantaneous modal amplitudes and accumulated convective history.
II.3 Permutation-Based Polynomial Construction
The key insight is that conserved quantities can be constructed systematically using permutations of the flow components as heuristic generators of polynomial candidates. We adopt a compact notation for permutations where we identify
$1\leftrightarrow x,\quad 2\leftrightarrow y,\quad 3\leftrightarrow z,\quad 4\leftrightarrow u,$ (7)
and write permutations as four-digit strings. For example, the permutation 1234 corresponds to the identity ordering $(x,y,z,u)$, while 2134 corresponds to $(y,x,z,u)$. When referring to permutation classes, we use the notation nabc where $n\in\{1,2,3,4\}$ is the fixed first index and $a,b,c$ denote the remaining indices in any order.
The inverse construction method. Rather than assuming that the gradient of a conserved quantity equals a permutation of the flow (which would violate Schwarz integrability conditions for most permutations), we employ the permutation as a generator to synthesize a polynomial candidate. Given a permutation $\pi$, we construct a candidate scalar $C$ by summing partial primitives:
$C(x,y,z,u)=\sum_{i=1}^{4}\int\pi(\mathbf{f})_{i}\,dq_{i},$ (8)
where $\pi(\mathbf{f})_{i}$ denotes the $i$-th component of the permuted flow vector and $\mathbf{q}=(x,y,z,u)$. This summation consolidates the contributions from each variable into a coherent algebraic structure, typically yielding a polynomial in $(x,y,z)$ plus terms linear in $u$.
The resulting candidate $C(x,y,z,u)$ generically takes the form
$C=P(x,y,z)+p(x,y,z)\cdot u,$ (9)
where $P$ is a polynomial and $p$ is a polynomial factor coupling the auxiliary variable to the physical coordinates.
Derivation of auxiliary dynamics. With $C$ constructed, we impose the conservation condition $\dot{C}=0$ and deduce the required auxiliary evolution. The conservation condition yields:
$\frac{\partial C}{\partial x}\dot{x}+\frac{\partial C}{\partial y}\dot{y}+\frac{\partial C}{\partial z}\dot{z}+\frac{\partial C}{\partial u}\dot{u}=0.$ (10)
Solving for $\dot{u}$:
$\dot{u}=-\frac{\frac{\partial C}{\partial x}\dot{x}+\frac{\partial C}{\partial y}\dot{y}+\frac{\partial C}{\partial z}\dot{z}}{\frac{\partial C}{\partial u}}=-\frac{\dot{P}+\dot{p}\cdot u}{p},$ (11)
where we used $\partial C/\partial u=p$. This equation reveals that $\dot{u}$ depends on $u$ itself through the term $(\dot{p}/p)u$. However, when we pass to the regularized variable $v=p\cdot u$, this dependence cancels exactly (Sec. III.2).
The systematic procedure thus consists of:
  1. 1.
    Enumerate all 24 permutations $\pi$ of the indices $\{1,2,3,4\}$.
  2. 2.
    For each permutation, construct the candidate $C$ via Eq. (8).
  3. 3.
    Derive the auxiliary evolution $\dot{u}$ from the conservation condition.
  4. 4.
    Check whether a valid (non-trivial, globally defined) invariant results.
This exhaustive analysis reveals that exactly 18 permutations yield valid invariants ($K_{1}$$K_{18}$), while 6 permutations yield null results. The null permutations are precisely those beginning with the index 4 (the auxiliary variable $u$), for which the integrability conditions fail at the first step (Theorem 1).
Classification by singularity structure. The polynomial $p(x,y,z)$ appearing in the denominator of Eq. (11) determines the singularity structure of the auxiliary evolution. Remarkably, the 18 valid permutations partition into exactly three families based on this polynomial:
  • Permutations 1abc yield p=yxp=y-x (Class I)
  • Permutations 2abc yield $p=y+x(z-\rho)$ (Class II)
  • Permutations 3abc yield $p=xy-\beta z$ (Class III)
This classification by regularization polynomial, rather than by permutation index alone, is the fundamental organizing principle. The correspondence between leading permutation index and class holds for the uniform-sign family explored here, but more general ansätze are classified solely by their regularization polynomial (see Sec. VIII.4).
It should be emphasized that the permutation-based construction provides sufficient conditions for the existence of conserved quantities, not necessary ones. There may exist history-dependent invariants of the Lorenz system that do not arise from any permutation ansatz. The present systematic exploration exhausts the permutation-based family but does not preclude the existence of structurally different invariants constructed by other methods.
II.4 Sign Convention
The invariants $K_{n}$ obtained through direct integration possess an intrinsic global gauge symmetry: if $K$ is a conserved quantity, then K-K is equally conserved. This $\mathbb{Z}_{2}$ symmetry $K\to-K$ reflects the mathematical arbitrariness in the overall sign of any constant of motion, a freedom absent in Noether-derived quantities, which inherit their signs from the underlying Lagrangian structure.
Throughout this work, we present the invariants in their canonical algorithmic form, the expressions that emerge directly from the integration procedure without additional sign manipulation. This choice ensures reproducibility and transparency of the constructive method. For physical applications, one may adopt an alternative positive-definite convention by applying global sign inversion $K_{n}\to-K_{n}$ to individual invariants such that dominant quadratic terms ($y^{2}$, $z^{2}$, $x^{2}$, etc.) appear with positive coefficients.
III Complete Family of Invariants
III.1 Organization by Permutation Class
The systematic exploration reveals an organizing principle: the first index in the permutation determines the regularization class. This leads to a natural partition:
  • Class I ($K_{1}$$K_{6}$): Permutations 1abc (first coordinate is $x$)
  • Class II ($K_{7}$$K_{12}$): Permutations 2abc (first coordinate is $y$)
  • Class III ($K_{13}$$K_{18}$): Permutations 3abc (first coordinate is $z$)
  • Null Class: Permutations 4abc (first coordinate is $u$)
The null class consists of permutations where the auxiliary variable $u$ occupies the first position. As we shall prove, these permutations cannot yield valid invariants because the resulting consistency equations admit only trivial solutions.
III.2 Regularization Polynomials and Invariant-Specific Auxiliary Variables
Each class is characterized by a distinct regularization polynomial $p(x,y,z)$ that eliminates singularities. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$ and regularized variable $v_{n}(t)$. Notation: Throughout this work, lowercase $p$ denotes the regularization polynomial (which multiplies the auxiliary variable), while uppercase $P$ denotes the polynomial part of the invariant (the non-auxiliary component of $K_{n}=P_{n}+v_{n}$).
Class I: $p_{\mathrm{I}}(x,y,z)=y-x,$
$v_{n}^{\mathrm{I}}=\sigma(y-x)u_{n},\quad n\in\{1,\ldots,6\}.$ (12)
Class II: $p_{\mathrm{II}}(x,y,z)=y+x(z-\rho),$
$v_{n}^{\mathrm{II}}=[y+x(z-\rho)]u_{n},\quad n\in\{7,\ldots,12\}.$ (13)
Class III: $p_{\mathrm{III}}(x,y,z)=xy-\beta z,$
$v_{n}^{\mathrm{III}}=(\beta z-xy)u_{n},\quad n\in\{13,\ldots,18\}.$ (14)
Note that $v_{n}^{\mathrm{III}}=-p_{\mathrm{III}}u_{n}$; this sign convention is adopted so that the polynomial parts $P_{n}$ of the Class III invariants have leading coefficients consistent with the other classes.
The subscript $n$ on $u_{n}$ and $v_{n}$ is essential: each invariant defines its own auxiliary variable. While all invariants within a class share the same regularization polynomial $p$, they differ in their polynomial parts $P_{n}(x,y,z)$, which determines the evolution equation for each auxiliary variable:
$\dot{v}_{n}=Q_{n}(x,y,z)\equiv-\frac{dP_{n}}{dt}.$ (15)
Since $P_{n}\neq P_{m}$ for $n\neq m$, we have $Q_{n}\neq Q_{m}$, and consequently $u_{n}(t)\neq u_{m}(t)$ along any trajectory.
These polynomials possess clear physical interpretations within the context of the Lorenz equations. The Class I polynomial $p_{\mathrm{I}}=y-x$ vanishes on the manifold where convective velocity equals horizontal temperature difference. The Class II polynomial $p_{\mathrm{II}}=y+x(z-\rho)$ is precisely the right-hand side of Eq. (2) with opposite sign, vanishing at the $y$-nullcline where $\dot{y}=0$. The Class III polynomial $p_{\mathrm{III}}=xy-\beta z$ is the right-hand side of Eq. (3), vanishing at the $z$-nullcline where $\dot{z}=0$. This connection to the nullcline structure of the Lorenz flow establishes that the regularization classes encode geometric features of the dynamics.
An important structural feature distinguishes these invariants from standard stochastic treatments: the auxiliary variable $u_{n}(t)$ enters multiplicatively through the coupling $p(x,y,z)\cdot u_{n}$, rather than as an additive noise term. In the Mori-Zwanzig picture, this corresponds to a state-dependent modulation of the memory contribution: the memory kernel’s influence on the dynamics is “gated” by the instantaneous values of the physical coordinates. This multiplicative structure has important consequences: when the polynomial $p(x,y,z)$ vanishes (on nullclines), the memory’s contribution to the invariant is momentarily suppressed, regardless of the accumulated history encoded in $u_{n}$. This gating mechanism may explain the topological robustness observed in the statistical analysis (Sec. VII), as the memory coupling is automatically reduced precisely where the flow undergoes critical transitions.
III.3 Structural Origin of the Regularized Evolution Equation
A key technical point deserves explicit attention: why does the evolution equation for the regularized variable $v_{n}$ emerge as a function of $(x,y,z)$ alone, independent of $v_{n}$ itself? This property is not assumed but follows structurally from the construction.
Consider an invariant of the general form $K_{n}=P_{n}(x,y,z)+p(x,y,z)\cdot u_{n}$. The conservation requirement $dK_{n}/dt=0$ yields:
$\frac{dP_{n}}{dt}+\frac{dp}{dt}u_{n}+p\dot{u}_{n}=0.$ (16)
Solving for $\dot{u}_{n}$ (assuming $p(x,y,z)\neq 0$; the behavior at nullclines where $p$ vanishes is addressed in Sec. III.4):
$\dot{u}_{n}=-\frac{1}{p}\left(\frac{dP_{n}}{dt}+\frac{dp}{dt}u_{n}\right)=-\frac{\dot{P}_{n}}{p}-\frac{\dot{p}}{p}u_{n}.$ (17)
The original variable $u_{n}$ has $u_{n}$-dependent dynamics. The term $(\dot{p}/p)u_{n}$ in Eq. (17) shows that $\dot{u}_{n}$ depends on $u_{n}$ itself. This means $u_{n}(t)$ satisfies a linear ODE with time-dependent coefficients, not a simple integral $u=\int f\,dt$. The $u_{n}$-dependent term is numerically stiff near nullclines where $p\to 0$.
The regularized variable $v_{n}$ has $v_{n}$-independent dynamics. Defining $v_{n}=p\cdot u_{n}$ and differentiating:
$\dot{v}_{n}=\dot{p}\cdot u_{n}+p\cdot\dot{u}_{n}=\dot{p}\cdot u_{n}+p\left(-\frac{\dot{P}_{n}}{p}-\frac{\dot{p}}{p}u_{n}\right)=-\dot{P}_{n}.$ (18)
The $u_{n}$-dependent terms cancel exactly, yielding $\dot{v}_{n}=Q_{n}(x,y,z)\equiv-\dot{P}_{n}$, which depends only on the physical coordinates. This algebraic cancellation is guaranteed by the structure of the invariant $K_{n}=P_{n}+v_{n}$: differentiating both sides of $K_{n}=c_{n}$ immediately gives $\dot{v}_{n}=-\dot{P}_{n}$ without reference to the intermediate variable $u_{n}$.
This distinction is essential: while the original auxiliary variable $u_{n}$ satisfies a stiff linear ODE, the regularized variable $v_{n}$ satisfies a simple integral $v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}\,ds$. The regularized formulation is therefore preferred for both theoretical analysis and numerical implementation.
Sign convention for Class III. The derivation above assumes $v=p\cdot u$. For Class III, we adopt the convention $v_{n}^{\mathrm{III}}=-p_{\mathrm{III}}\cdot u_{n}$ (Eq. 14). This sign flip does not affect the fundamental result $\dot{v}_{n}=-\dot{P}_{n}$, which follows directly from differentiating $K_{n}=P_{n}+v_{n}=c_{n}$ regardless of how $v_{n}$ is defined in terms of $u_{n}$. However, the evolution equation for the original auxiliary variable $u_{n}$ differs between classes. Since all analysis in this work employs the regularized variables $v_{n}$ exclusively, this sign convention affects only the formal relationship between $u_{n}$ and $v_{n}$, not the invariant calculations themselves.
Explicit verification for $K_{13}$. To confirm sign consistency, we verify directly that $K_{13}=P_{13}+v_{13}$ is conserved. With $P_{13}=y^{2}/2+xyz-xy\rho+x^{2}\sigma/2-xy\sigma$ and $v_{13}^{\mathrm{III}}=(\beta z-xy)u_{13}$, the invariant is:
$K_{13}=\frac{y^{2}}{2}+xyz-xy\rho+\frac{x^{2}\sigma}{2}-xy\sigma+v_{13}.$ (19)
Differentiating: $\dot{K}_{13}=\dot{P}_{13}+\dot{v}_{13}$. Setting $\dot{v}_{13}=-\dot{P}_{13}$ yields $\dot{K}_{13}=0$ as required. The evolution function $Q_{13}=\dot{v}_{13}=-\dot{P}_{13}$ exhibits large-amplitude spikes at separatrix crossings (as derived in Sec. VII.4), consistent with the intermittent behavior observed in Figure 2(c) and the elevated skewness magnitude $|S_{\mathrm{III}}|\approx 2.7$.
The cancellation fails for the null class (permutations 4abc) precisely because placing $u$ in the leading position violates the integrability conditions (Schwarz’s theorem) before the regularization step can be applied, as demonstrated in Theorem 1.
III.4 Singularity Structure and the Distinction Between $u_{n}$ and $v_{n}$
The regularization polynomials $p_{\mathrm{I}}$, $p_{\mathrm{II}}$, and $p_{\mathrm{III}}$ vanish on nullclines of the Lorenz flow. This has different consequences for $u_{n}$ and $v_{n}$.
The regularized variable $v_{n}$ is everywhere smooth and bounded. From conservation $K_{n}=P_{n}+v_{n}=c_{n}$:
$v_{n}(t)=c_{n}-P_{n}(x(t),y(t),z(t)).$ (20)
Since $P_{n}$ is polynomial and the attractor is compact, $v_{n}$ inherits bounded, smooth behavior. The evolution equation $\dot{v}_{n}=Q_{n}(x,y,z)$ involves only polynomial functions, ensuring $v_{n}$ varies continuously across nullclines.
The original auxiliary variable $u_{n}$ diverges at nullcline crossings. The relationship $v_{n}=p_{n}\cdot u_{n}$ implies
$u_{n}(t)=\frac{v_{n}(t)}{p_{n}(x(t),y(t),z(t))}=\frac{c_{n}-P_{n}(x,y,z)}{p_{n}(x,y,z)}.$ (21)
When the trajectory crosses a nullcline where $p_{n}=0$, $u_{n}\to\pm\infty$.
This divergence is a coordinate singularity, analogous to the singularity at $r=0$ in polar coordinates, not a physical singularity. The invariant $K_{n}=P_{n}+p_{n}u_{n}=P_{n}+v_{n}$ remains well-defined:
$\lim_{p_{n}\to 0}K_{n}=\lim_{p_{n}\to 0}\left[P_{n}+p_{n}\cdot\frac{c_{n}-P_{n}}{p_{n}}\right]=c_{n}.$ (22)
Proper formulation. The rigorous formulation employs regularized variables:
$K_{n}=P_{n}(x,y,z)+v_{n},\quad\dot{v}_{n}=Q_{n}(x,y,z).$ (23)
This representation is singularity-free throughout phase space.
III.5 Class I Invariants: Permutations 1abc
The six Class I invariants have form $K_{n}=P_{n}(x,y,z)+v_{n}^{\mathrm{I}}$ where $v_{n}^{\mathrm{I}}=\sigma(y-x)u_{n}$:
$K_{1}\text{(1234)}$ $=xy-\frac{xy^{2}}{2}+\frac{x^{2}z}{2}+yz\beta-\frac{x^{2}\rho}{2}+v_{1}^{\mathrm{I}},$ (24)
$K_{2}\text{(1243)}$ $=-\frac{x^{2}y}{2}+\frac{y^{2}}{2}+xyz+xz\beta-xy\rho+v_{2}^{\mathrm{I}},$ (25)
$K_{3}\text{(1324)}$ $=-\frac{xy^{2}}{2}+yz+\frac{xz^{2}}{2}+yz\beta-xz\rho+v_{3}^{\mathrm{I}},$ (26)
$K_{4}\text{(1342)}$ $=-\frac{x^{2}y}{2}+yz+\frac{xz^{2}}{2}+xz\beta-xz\rho+v_{4}^{\mathrm{I}},$ (27)
$K_{5}\text{(1423)}$ $=\frac{y^{2}}{2}+\frac{z^{2}\beta}{2}-xy\rho+v_{5}^{\mathrm{I}},$ (28)
$K_{6}\text{(1432)}$ $=xy+\frac{x^{2}z}{2}-xyz+\frac{z^{2}\beta}{2}-\frac{x^{2}\rho}{2}+v_{6}^{\mathrm{I}}.$ (29)
III.6 Class II Invariants: Permutations 2abc
The six Class II invariants have form $K_{n}=P_{n}(x,y,z)+v_{n}^{\mathrm{II}}$ where $v_{n}^{\mathrm{II}}=[y+x(z-\rho)]u_{n}$:
$K_{7}\text{(2134)}$ $=-\frac{xy^{2}}{2}+yz\beta+\frac{x^{2}\sigma}{2}-xy\sigma+v_{7}^{\mathrm{II}},$ (30)
$K_{8}\text{(2143)}$ $=-\frac{x^{2}y}{2}+xz\beta+xy\sigma-\frac{y^{2}\sigma}{2}+v_{8}^{\mathrm{II}},$ (31)
$K_{9}\text{(2314)}$ $=-\frac{xy^{2}}{2}+yz\beta+xz\sigma-yz\sigma+v_{9}^{\mathrm{II}},$ (32)
$K_{10}\text{(2341)}$ $=-\frac{x^{2}y}{2}+xz\beta+xz\sigma-yz\sigma+v_{10}^{\mathrm{II}},$ (33)
$K_{11}\text{(2413)}$ $=-xyz+\frac{z^{2}\beta}{2}+xy\sigma-\frac{y^{2}\sigma}{2}+v_{11}^{\mathrm{II}},$ (34)
$K_{12}\text{(2431)}$ $=-xyz+\frac{z^{2}\beta}{2}+\frac{x^{2}\sigma}{2}-xy\sigma+v_{12}^{\mathrm{II}}.$ (35)
III.7 Class III Invariants: Permutations 3abc
The six Class III invariants have form $K_{n}=P_{n}(x,y,z)+v_{n}^{\mathrm{III}}$ where $v_{n}^{\mathrm{III}}=(\beta z-xy)u_{n}$:
$K_{13}\text{(3124)}$ $=\frac{y^{2}}{2}+xyz-xy\rho+\frac{x^{2}\sigma}{2}-xy\sigma+v_{13}^{\mathrm{III}},$ (36)
$K_{14}\text{(3142)}$ $=xy+\frac{x^{2}z}{2}-\frac{x^{2}\rho}{2}+xy\sigma-\frac{y^{2}\sigma}{2}+v_{14}^{\mathrm{III}},$ (37)
$K_{15}\text{(3214)}$ $=\frac{y^{2}}{2}+xyz-xy\rho+xz\sigma-yz\sigma+v_{15}^{\mathrm{III}},$ (38)
$K_{16}\text{(3241)}$ $=xy+\frac{x^{2}z}{2}-\frac{x^{2}\rho}{2}+xz\sigma-yz\sigma+v_{16}^{\mathrm{III}},$ (39)
$K_{17}\text{(3412)}$ $=yz+\frac{xz^{2}}{2}-xz\rho+xy\sigma-\frac{y^{2}\sigma}{2}+v_{17}^{\mathrm{III}},$ (40)
$K_{18}\text{(3421)}$ $=yz+\frac{xz^{2}}{2}-xz\rho+\frac{x^{2}\sigma}{2}-xy\sigma+v_{18}^{\mathrm{III}}.$ (41)
III.8 Null Class: Permutations 4abc
The six permutations beginning with 4 yield no valid invariants.
Theorem 1 (Null Permutations).
Permutations 4abc yield no valid history-dependent invariants.
Proof.
We demonstrate for permutation 4123, corresponding to $(u,x,y,z)$. The orthogonality ansatz assigns
$\nabla C=(\dot{u},\dot{x},\dot{y},\dot{z}),$ (42)
yielding:
$\partial_{x}C$ $=\dot{u}=f(x,y,z,u),$ (43)
$\partial_{y}C$ $=\dot{x}=\sigma(y-x),$ (44)
$\partial_{z}C$ $=\dot{y}=x(\rho-z)-y,$ (45)
$\partial_{u}C$ $=\dot{z}=xy-\beta z.$ (46)
Consider Schwarz’s theorem for mixed partial derivatives. From Eqs. (44) and (46):
$\partial_{u}(\partial_{y}C)$ $=\partial_{u}[\sigma(y-x)]=0,$ (47)
$\partial_{y}(\partial_{u}C)$ $=\partial_{y}[xy-\beta z]=x.$ (48)
Equality $\partial_{yu}C=\partial_{uy}C$ requires $0=x$. This constraint restricts $C$ to the hyperplane $x=0$, which is not invariant under the Lorenz flow. No smooth invariant satisfying ansatz (42) exists throughout phase space.
For valid classes (1abc, 2abc, 3abc), the ansatz assigns $\partial_{u}C$ to $\dot{u}$, which is determined self-consistently. The flexibility to choose $\dot{u}$ allows the integrability conditions to be satisfied.
For null permutations 4abc, $\partial_{u}C$ is assigned to one of $\dot{x}$, $\dot{y}$, or $\dot{z}$, whose functional forms are prescribed by Eqs. (1)–(3). This eliminates the flexibility needed for consistency: the system becomes overdetermined.
Analogous contradictions arise for all permutations 4abc. Permutations 4213 and 4231 yield the constraint $0=y$; permutations 4312 and 4321 yield $0=\sigma(y-x)$; permutation 4132 yields $0=x$. Each constraint defines a lower-dimensional manifold not invariant under the flow. ∎
This result reflects violation of Schwarz integrability conditions when the auxiliary variable occupies the leading position. The ansatz generates an overdetermined system with no smooth solution throughout phase space.
Table 1: Master table of permutations: correspondence between invariant index $K_{n}$ and generating permutation $\pi$. Indices 1–4 map to coordinates $x,y,z,u$ respectively.
Class I Class II Class III
$K_{n}$ $\pi$ $K_{n}$ $\pi$ $K_{n}$ $\pi$
$K_{1}$ 1234 $K_{7}$ 2134 $K_{13}$ 3124
$K_{2}$ 1243 $K_{8}$ 2143 $K_{14}$ 3142
$K_{3}$ 1324 $K_{9}$ 2314 $K_{15}$ 3214
$K_{4}$ 1342 $K_{10}$ 2341 $K_{16}$ 3241
$K_{5}$ 1423 $K_{11}$ 2413 $K_{17}$ 3412
$K_{6}$ 1432 $K_{12}$ 2431 $K_{18}$ 3421
Null Class: Permutations 4123, 4132, 4213, 4231, 4312, 4321 yield no valid invariants due to Schwarz integrability constraints.
[…]
IV Functional Independence and Auxiliary Variable Structure
IV.1 Boundedness of Regularized Variables
A consequence of the conservation laws is that regularized variables $v_{n}(t)$ remain bounded for all time, despite their coupling to the trajectory history through $\dot{v}_{n}=Q_{n}(x,y,z)$.
From conservation $K_{n}=P_{n}(x,y,z)+v_{n}=c_{n}$:
$v_{n}(t)=c_{n}-P_{n}(x(t),y(t),z(t)).$ (49)
Since $P_{n}$ is polynomial and the Lorenz attractor is compact, $v_{n}(t)$ is bounded. The fluctuation range is determined by the variation of $P_{n}$ over the attractor.
Distinction from $u_{n}$. The original auxiliary variable $u_{n}(t)=v_{n}(t)/p_{n}(x,y,z)$ diverges at nullcline crossings (Sec. III.4). This coordinate singularity does not affect conservation. We work exclusively with regularized variables $v_{n}$.
Zero time-average. Boundedness implies that $Q_{n}=\dot{v}_{n}$ has zero time-average with respect to the ergodic measure:
$\langle Q_{n}\rangle_{\mathrm{SRB}}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}Q_{n}(x(t),y(t),z(t))\,dt=0.$ (50)
A nonzero average would imply unbounded growth of $v_{n}$, contradicting Eq. (49).
Numerical integration at standard parameters over $T=5000$ confirms these predictions with high precision.
IV.2 Structure of Pairwise Differences
Boundedness extends to differences of regularized variables. For invariants $K_{n}$ and $K_{m}$:
$v_{n}-v_{m}=(c_{n}-P_{n})-(c_{m}-P_{m})=(c_{n}-c_{m})-(P_{n}-P_{m}).$ (51)
Since $P_{n}-P_{m}$ is polynomial and bounded on the attractor, $v_{n}-v_{m}$ is bounded.
For invariants from different regularization classes, this algebraic synchronization reflects a structural property: although regularized variables follow different evolution equations, they remain coupled through the shared trajectory. The conservation laws enforce a coherence that prevents unbounded divergence.
IV.3 Physical Interpretation: Bounded History Accumulation
Each invariant $K_{n}$ defines a “measure of history” through $v_{n}(t)$, but this history does not accumulate indefinitely. It fluctuates within a bounded range determined by the attractor’s geometry.
From Eq. (49), the regularized variable at any instant is determined by the current state $(x,y,z)$ and the conserved value $c_{n}$. The evolution equation $v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}\,ds$ and the algebraic representation $v_{n}=c_{n}-P_{n}$ are equivalent, but they emphasize different aspects. The former emphasizes history-dependence; the latter reveals that “history” is encoded in the instantaneous state.
This equivalence requires clarification to avoid logical ambiguity. While $v_{n}(t)=c_{n}-P_{n}(x,y,z)$ appears to render $v_{n}$ a function of the instantaneous state alone, this representation presupposes knowledge of the conserved value $c_{n}$. The quantity $c_{n}$ is determined by the initial condition: $c_{n}=P_{n}(x_{0},y_{0},z_{0})$ when $v_{n}(0)=0$. Crucially, $c_{n}$ cannot be recovered from the current state $(x(t),y(t),z(t))$ without knowledge of the initial state or, equivalently, the path integral $\int_{0}^{t}Q_{n}\,ds$. The “non-local” character of the constraint arises precisely because the instantaneous state does not determine the invariant value; the initial condition is required. In this sense, the constraint encodes information about the trajectory’s origin that is not contained in its current position.
Different invariants assign different “variations” to the same trajectory segment. When the trajectory passes through regions where $P_{n}$ changes rapidly, $v_{n}$ adjusts correspondingly. The rate of adjustment is encoded in $Q_{n}=\dot{v}_{n}$. The statistical analysis (Sec. VII) shows that these variations differ across classes: Class III invariants incur large-amplitude spikes in $Q_{\mathrm{III}}$ at lobe-switching events, while Class I invariants accumulate more smoothly.
The fluctuations of $v_{n}(t)$ trace a characteristic signature that encodes the trajectory’s relationship to the nullcline structure. The class structure partitions the invariants by dynamical sensitivity: Class III invariants respond to discrete topological transitions, while Class I invariants respond to continuous dynamics.
IV.4 Constraints on Trajectory Histories
Theorem 2 (Independence of History Constraints).
For any selection of three invariants $\{K_{\mathrm{I}},K_{\mathrm{II}},K_{\mathrm{III}}\}$ with one from each regularization class, the corresponding constraints are functionally independent:
$K_{\mathrm{I}}$ $=P_{\mathrm{I}}(x,y,z)+v_{\mathrm{I}}=c_{1},$ (52)
$K_{\mathrm{II}}$ $=P_{\mathrm{II}}(x,y,z)+v_{\mathrm{II}}=c_{2},$ (53)
$K_{\mathrm{III}}$ $=P_{\mathrm{III}}(x,y,z)+v_{\mathrm{III}}=c_{3},$ (54)
provide three independent conditions that, together with the initial state $(x_{0},y_{0},z_{0})$, characterize a trajectory up to the discrete $\mathbb{Z}_{2}$ symmetry.
Proof.
Independence follows from the algebraic independence of the regularization polynomials. At $t=0$, with $v_{n}(0)=0$ (assuming $u_{n}(0)=0$), the invariant values are:
$c_{n}=P_{n}(x_{0},y_{0},z_{0}).$ (55)
Since $P_{\mathrm{I}}$, $P_{\mathrm{II}}$, and $P_{\mathrm{III}}$ are algebraically distinct polynomials, the map $(x_{0},y_{0},z_{0})\mapsto(c_{1},c_{2},c_{3})$ has full rank at generic points. Given $(c_{1},c_{2},c_{3})$, the polynomial equations $P_{n}(x,y,z)=c_{n}$ determine a discrete set of initial conditions.
For the canonical Triad $\{K_{5},K_{11},K_{13}\}$, the $3\times 3$ Jacobian matrix $\tilde{\mathbf{J}}$ of polynomial parts:
$\tilde{\mathbf{J}}=\begin{pmatrix}\partial_{x}P_{5}&\partial_{y}P_{5}&\partial_{z}P_{5}\\ \partial_{x}P_{11}&\partial_{y}P_{11}&\partial_{z}P_{11}\\ \partial_{x}P_{13}&\partial_{y}P_{13}&\partial_{z}P_{13}\end{pmatrix}$ (56)
has leading-order determinant for large $z$:
$\det(\tilde{\mathbf{J}})=\beta\sigma z^{2}(y-x)(y-x\rho)+\text{lower-order terms}.$ (57)
The full determinant is degree 5 in $(x,y,z)$; the zero locus is a lower-dimensional algebraic variety. ∎
Degeneracy locus. The determinant (57) vanishes on:
  1. 1.
    $z=0$: the horizontal plane through the origin;
  2. 2.
    $y=x$: the diagonal plane (Class I nullcline);
  3. 3.
    $y=x\rho$: a plane through the origin with slope $\rho$.
These loci have measure zero with respect to both Lebesgue measure and the SRB measure on the attractor. The fingerprint $(c_{1},c_{2},c_{3})$ provides a unique identifier at generic points.
IV.5 The Triad Construction
Selecting one invariant from each regularization class,
$\mathcal{T}=\{K_{\mathrm{I}},K_{\mathrm{II}},K_{\mathrm{III}}\},$ (58)
provides a canonical set of three independent history constraints. The specific choice within each class affects only the numerical values, not the structural information.
Each Triad defines a “fingerprint” $(c_{1},c_{2},c_{3})$ that can be computed instantaneously from $(x_{0},y_{0},z_{0})$ via Eq. (55), since $v_{n}(0)=0$. The canonical Triad $\{K_{5},K_{11},K_{13}\}$ consists of even invariants under $\mathbb{Z}_{2}$, so the fingerprint identifies trajectories up to discrete symmetry.
V Geometric Interpretation
V.1 Invariants as Constraints on Trajectory Histories
The history-dependent invariants differ qualitatively from classical conservation laws. In Hamiltonian mechanics, $H(q,p)=E$ defines a hypersurface in instantaneous phase space; trajectories are confined to this surface for all time. The constraint is local in time.
The invariants $K_{n}=P_{n}(x,y,z)+v_{n}=c_{n}$ define constraints of a different character. The regularized variable $v_{n}(t)$ is a functional of the trajectory history:
$v_{n}[\gamma]=\int_{\gamma}Q_{n}(x,y,z)\,dt,$ (59)
where $\gamma$ denotes the trajectory from $t=0$ to the current time. The constraint $K_{n}=c_{n}$ relates the instantaneous state $(x,y,z)$ to the accumulated history $v_{n}$; it is non-local in time.
Different invariants require different regularized variables because each imposes a distinct relationship between configuration and history. Invariants from different regularization classes “observe” the trajectory through different lenses: Class I through the yxy-x nullcline structure, Class III through the $z$-nullcline and its connection to lobe-switching events.
V.2 Visualization of Constraint Surfaces
One can fix the regularized variable $v_{n}$ at a particular value and plot the resulting surface in $(x,y,z)$ space. Figure 1 displays such surfaces for the canonical Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$ along an unstable periodic orbit (UPO).
Several caveats apply:
  1. 1.
    The three surfaces correspond to different regularized variables $(v_{5},v_{11},v_{13})$, not a shared coordinate.
  2. 2.
    Along the UPO, regularized variables take specific values determined by integration over the orbit.
  3. 3.
    The surfaces intersect along the UPO because the invariant values were computed from this orbit’s initial conditions.
For any trajectory, the three constraints $K_{n}=c_{n}$ are satisfied simultaneously. Each constraint operates in its own extended space $(x,y,z,v_{n})$.
V.3 Resolution of the Structure-Chaos Relationship
The invariants illuminate the relationship between structure and chaos in the Lorenz system.
Structure: Each trajectory is characterized by its Triad fingerprint $(c_{1},c_{2},c_{3})$, determined instantaneously from initial conditions. The canonical Triad consists of even invariants, so the fingerprint identifies trajectories up to $\mathbb{Z}_{2}$ symmetry. Different trajectories (generically) have different fingerprints.
Chaos: Chaotic dynamics arise from sensitivity to perturbations in the fingerprint. Infinitesimally different initial conditions produce infinitesimally different fingerprints. While each fingerprint uniquely determines a trajectory, the mapping from fingerprint to long-time behavior inherits the Lyapunov instability of the underlying flow.
This resolves the apparent tension between “chaos” and “structure”: the invariants provide exact organization (a unique fingerprint per trajectory), while chaos persists as the exponential divergence of trajectories with nearby fingerprints.
V.4 Computational Complexity
The history-dependent constraints do not simplify the prediction of future states. Determining $(x(t),y(t),z(t))$ from $(x_{0},y_{0},z_{0})$ still requires numerical integration of the Lorenz equations. The regularized variable $v_{n}(t)=\int_{0}^{t}Q_{n}\,ds$ requires knowledge of the complete trajectory up to time $t$.
The computational complexity in the invariant formulation equals that of the original Lorenz system. The constraints provide structural insight and classification tools (Triad fingerprints), but do not reduce the dynamics to quadratures as classical integrability would.
V.5 Connection to Inertial Manifolds
The invariant structure resonates with, but differs from, the theory of inertial manifolds in dissipative systems [12]. An inertial manifold is a finite-dimensional, exponentially attracting, positively invariant smooth manifold that contains the global attractor.
The differences are significant:
  1. 1.
    Temporal validity: Inertial manifolds describe asymptotic behavior ($t\to\infty$). The invariant constraints hold exactly for all $t\geq 0$.
  2. 2.
    Mechanism: Inertial manifolds arise from spectral gaps in the linearized dynamics. Invariant constraints arise from the algebraic structure of the embedding.
  3. 3.
    Nature of constraint: Inertial manifolds are geometric objects in the instantaneous state space. Invariant constraints involve path functionals and are non-local in time.
V.6 Periodic Orbit Classification
Unstable periodic orbits (UPOs) form the “skeleton” of the strange attractor [13]. Their enumeration and classification are complicated by numerical instability and the lack of intrinsic invariant labels.
The Triad fingerprint provides a resolution. Each UPO corresponds to a specific triple $(c_{1},c_{2},c_{3})$. This serves as an intrinsic fingerprint that is independent of how the orbit is parameterized or which point is chosen as the initial condition. Two orbits with the same fingerprint are identical (up to $\mathbb{Z}_{2}$ symmetry); distinct orbits have different fingerprints.
This transforms UPO enumeration from a numerical search problem into a geometric cataloguing problem.
V.7 Poincaré Section Construction
The visualization of constraint surface intersections (Fig. 1b) requires a carefully constructed Poincaré section. A standard canonical section (e.g., $z=\mathrm{const}$) produces tangent or nearly degenerate curve intersections that obscure the independence of the three constraints. We therefore employ an optimized generalized section.
Reference point selection. The section plane $\Pi$ is centered at a point $\vec{P}_{\mathrm{opt}}$ on the unstable periodic orbit. Starting from the UPO initial condition $(x_{0},y_{0},z_{0})=(-1.31956,-0.38627,22.60542)$ with period $T\approx 1.55865$, the cutting point is defined at time
$t_{\mathrm{cut}}=T/4+0.0587\approx 0.4484,$ (60)
yielding $\vec{P}_{\mathrm{opt}}=\vec{X}(t_{\mathrm{cut}})$. This point serves as the origin of the local coordinate system on the section.
Normal vector optimization. The section plane is defined by a normal vector $\vec{n}(\theta,\phi)$ parameterized in spherical coordinates. The optimal orientation is determined by maximizing the minimum angular separation between projected constraint curves. For each candidate normal satisfying the transversality condition $|\vec{n}\cdot\vec{v}|>0.1$ (where $\vec{v}$ is the flow velocity), the tangent vectors of the projected level curves are computed as $\vec{t}_{i}=\vec{n}\times\nabla K_{i}$. The objective function
$\mathcal{J}(\theta,\phi)=\min_{i\neq j}\angle(\vec{t}_{i},\vec{t}_{j})$ (61)
is maximized using differential evolution, ensuring that the three constraint curves intersect with maximum angular aperture.
Local coordinates. The two-dimensional section employs orthonormal coordinates $(\eta,\xi)$ defined on $\Pi$. The basis vectors are constructed as
[…]
$\vec{V}$ $=\mathrm{Normalize}(\vec{n}\times\vec{U}),$ (63)
where $\hat{k}$ is the unit vector along $z$ (or $\hat{j}$ if singular). Any point $(\eta,\xi)$ in the section maps to three-dimensional space via
$\vec{X}_{3D}=\vec{P}_{\mathrm{opt}}+\eta\,\vec{U}+\xi\,\vec{V}.$ (64)
The coordinates $\eta$ and $\xi$ represent Euclidean displacements on the optimized plane, ensuring undistorted visualization of the constraint geometry.
Figure 1: Visualization of history-dependent constraints in $(x,y,z)$ space. (a) An unstable periodic orbit (dark curve) at standard parameters with constraint surfaces for Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$: Class I surface (blue), Class II surface (green), Class III surface (red). The surfaces represent $K_{n}=c_{n}$ at fixed $v_{n}$ values evaluated along the UPO. (b) Optimized Poincaré section through $\vec{P}_{\mathrm{opt}}$ on the UPO. The constraint surfaces intersect the plane as curves: $K_{5}=c_{1}$ (blue), $K_{11}=c_{2}$ (green), $K_{13}=c_{3}$ (red). The curves meet at a single point (the UPO crossing) with maximized angular separation, demonstrating the geometric independence of the three constraints. Coordinates $(\eta,\xi)$ are Euclidean displacements on the optimized section plane.
VI Symmetry Analysis
VI.1 The Discrete $\mathbb{Z}_{2}$ Symmetry
The Lorenz system possesses the discrete symmetry
$\mathcal{S}:(x,y,z)\to(-x,-y,z),$ (65)
which leaves Eqs. (1)–(3) invariant.
The regularization polynomials transform under $\mathcal{S}$ as follows:
$p_{\mathrm{I}}$ $=y-x\to-p_{\mathrm{I}}\quad\text{(odd)},$ (66)
$p_{\mathrm{II}}$ $=y+x(z-\rho)\to-p_{\mathrm{II}}\quad\text{(odd)},$ (67)
$p_{\mathrm{III}}$ $=xy-\beta z\to p_{\mathrm{III}}\quad\text{(even)}.$ (68)
VI.2 Parity Classification of Invariants
The invariants can be classified by the transformation properties of their polynomial parts $P_{n}(x,y,z)$ under $\mathcal{S}$. The bilinear terms transform as:
$x^{2},y^{2},z^{2},xy,xyz\to\text{(even)},$
$xz,yz\to\text{(odd)}.$ (69)
Table 2 presents the complete parity classification. Six invariants ($K_{5}$, $K_{6}$, $K_{11}$, $K_{12}$, $K_{13}$, $K_{14}$) are purely even; four ($K_{3}$, $K_{4}$, $K_{9}$, $K_{10}$) are purely odd; eight possess mixed parity.
Table 2: Classification of invariants by $\mathbb{Z}_{2}$ parity of $P_{n}$. E = even; O = odd; M = mixed.
Class Invariants Parity Odd terms
I $K_{5}$, $K_{6}$ E
$K_{3}$, $K_{4}$ O yzyz, xzxz
$K_{1}$, $K_{2}$ M yzyz, xzxz
II $K_{11}$, $K_{12}$ E
$K_{9}$, $K_{10}$ O yzyz, xzxz
$K_{7}$, $K_{8}$ M yzyz, xzxz
III $K_{13}$, $K_{14}$ E
$K_{15}$$K_{18}$ M yzyz, xzxz
VI.3 Implications for Triad Classification
The canonical Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$ consists entirely of even invariants. Therefore
$P_{n}(\mathcal{S}(x,y,z))=P_{n}(x,y,z)$ (70)
for $n\in\{5,11,13\}$. Symmetric initial conditions produce identical fingerprints:
$(c_{1},c_{2},c_{3})\big|_{(x_{0},y_{0},z_{0})}=(c_{1},c_{2},c_{3})\big|_{(-x_{0},-y_{0},z_{0})}.$ (71)
The fingerprint therefore characterizes trajectories up to $\mathbb{Z}_{2}$ symmetry. Symmetric orbits share fingerprints, reflecting the fact that they are dynamically equivalent.
Parity-breaking Triads. To distinguish symmetric trajectories, one can employ an odd invariant. The Triad $\mathcal{T}^{\prime}=\{K_{3},K_{11},K_{13}\}$ includes the odd invariant $K_{3}$, yielding $(c_{1}^{\prime},c_{2},c_{3})|_{\mathcal{S}(\gamma)}=(-c_{1}^{\prime},c_{2},c_{3})|_{\gamma}$.
VI.4 Regularization Class and Sensitivity to Lobe Transitions
The Lorenz attractor possesses topological structure encoded in its template [6]. Periodic orbits can be classified by their symbolic dynamics: the sequence of left (L) and right (R) lobe visits.
The Class III polynomial parts $P_{\mathrm{III}}$ contain terms involving $xy$. Since the product $xy$ changes sign at lobe-switching events (when the trajectory crosses $x=0$), $\dot{P}_{\mathrm{III}}$ varies rapidly at these transitions.
At the separatrix, $\dot{x}=\sigma(y-x)$, so at $x=0$: $\dot{x}=\sigma y$. The product $xy$ evolves from negative (before crossing) through zero (at crossing) to positive (after crossing) during a left-to-right transition. The derivative $d(xy)/dt|_{x=0}=\sigma y^{2}>0$ confirms monotonic increase through crossing.
Class III invariants are therefore sensitive to separatrix position. The evolution function $Q_{\mathrm{III}}=-\dot{P}_{\mathrm{III}}$ exhibits large magnitude at lobe crossings. Class I polynomial parts also contain $xy$ terms but include $z^{2}$ contributions that distribute variance across the entire trajectory, reducing the relative prominence of separatrix spikes (Sec. VII.4).
VII Numerical Verification and Statistical Analysis
VII.1 High-Precision Verification
We performed numerical integration using 30-digit extended precision arithmetic in Mathematica with the StiffnessSwitching method, which automatically adapts to handle stiff intervals during separatrix crossings.
For each invariant $K_{n}$, we verified:
  1. 1.
    Conservation: $|K_{n}(T)-K_{n}(0)|<10^{-25}$ over $T=5000$ time units.
  2. 2.
    Boundedness: $v_{n}(t)$ fluctuates within a finite range consistent with $v_{n}=c_{n}-P_{n}$.
  3. 3.
    Zero mean: $|\langle Q_{n}\rangle_{T}|<10^{-10}$ for $T>1000$.
Table 3 presents conservation errors for the canonical Triad after one UPO period. Errors of order $10^{-36}$$10^{-38}$ confirm conservation to machine precision.
Table 3: Conservation errors |K(T)K(0)||K(T)-K(0)| for Triad invariants after one UPO period ($T\approx 1.56$).
Invariant Class Error
$K_{5}$ I $3.45\times 10^{-36}$
$K_{11}$ II $9.35\times 10^{-38}$
$K_{13}$ III $3.68\times 10^{-38}$
VII.2 Statistical Analysis of Auxiliary Evolution Functions
We examine whether invariants from different regularization classes exhibit distinct dynamical signatures. For each invariant $K_{n}=P_{n}(x,y,z)+v_{n}$, the evolution function
$Q_{n}(x,y,z)\equiv\dot{v}_{n}=-\frac{dP_{n}}{dt}$ (72)
represents the instantaneous rate at which $v_{n}$ evolves to maintain conservation.
Pre-chaotic regime ($\rho=23$). Both classes yield statistically indistinguishable distributions (Fig. 2a). Kurtosis values $\kappa_{\mathrm{I}}\approx 36.8$ and $\kappa_{\mathrm{III}}\approx 37.0$ are effectively identical. This baseline establishes that there is no intrinsic bias in the Class III formulation.
Chaotic regime ($\rho=28$). A clear structural divergence emerges (Fig. 2b). Standardized distributions reveal different tail structures:
Differential intermittency. Class III exhibits higher kurtosis ($\kappa_{\mathrm{III}}\approx 15.1$) compared to Class I ($\kappa_{\mathrm{I}}\approx 7.9$), a 91% increase. The heavy tails indicate that Class III conservation is subject to sporadic large-amplitude events (“spikes”) punctuating quiescent intervals.
Asymmetry structure. Class III displays stronger negative skewness ($S_{\mathrm{III}}\approx-2.67$) versus Class I ($S_{\mathrm{I}}\approx-1.64$). Negative skewness indicates that extreme negative excursions are more intense than positive fluctuations. At separatrix crossings in the upper region of the attractor ($z>38$), $\dot{P}_{\mathrm{III}}>0$, yielding $Q_{\mathrm{III}}=-\dot{P}_{\mathrm{III}}<0$. The enhanced negative skewness reflects the concentration of large negative spikes at lobe switches.
Interpretation. The elevated kurtosis correlates with lobe-switching events. At separatrix crossings, $P_{\mathrm{III}}$ varies rapidly due to its $xy$-terms, forcing rapid adjustments in the Class III evolution function. Class I quantities lack this geometric sensitivity.
These dynamical signatures establish that the regularization class structure reflects genuine differences arising from polynomial structure, not algebraic convention.
Table 4: Statistical characterization of evolution functions $Q_{n}(t)$ for Class I ($K_{5}$) and Class III ($K_{13}$). Kurtosis $\kappa$ measures tail weight (Gaussian: $\kappa=3$); skewness $S$ measures asymmetry. Statistics computed from $N=9\times 10^{4}$ samples over $T=1000$ time units.
$\rho=23$ (pre-chaotic) $\rho=28$ (chaotic)
Statistic $Q_{\mathrm{I}}$ $Q_{\mathrm{III}}$ $Q_{\mathrm{I}}$ $Q_{\mathrm{III}}$
Kurtosis $\kappa$ 36.8 37.0 7.9 15.1
Skewness $S$ 0.33-0.33 0.38-0.38 1.64-1.64 2.67-2.67
Pre-chaotic versus chaotic kurtosis. A noteworthy feature of Table 4 is that the pre-chaotic regime ($\rho=23$) exhibits substantially higher kurtosis ($\kappa\approx 37$) than the fully chaotic regime ($\kappa_{\mathrm{III}}\approx 15$). This counterintuitive result reflects a fundamental difference in the statistical structure of the two dynamical regimes. At $\rho=23$, the system operates near the subcritical Hopf bifurcation where trajectories exhibit complex transient dynamics: long intervals of quiescence near local attractors or unstable periodic orbits are punctuated by abrupt escapes. These rare but intense bursts generate distributions with pronounced heavy tails, yielding elevated kurtosis. In contrast, the fully developed chaos at $\rho=28$ produces more ergodic and mixing dynamics, with fluctuations distributed more uniformly across time. The increased frequency of moderate-amplitude events dilutes the statistical prominence of extreme excursions, thereby reducing kurtosis. Additionally, Class I and Class III distributions are statistically indistinguishable at $\rho=23$ ($\kappa\approx 37$ for both), while they diverge significantly at $\rho=28$ ($\kappa_{\mathrm{I}}=7.9$ vs. $\kappa_{\mathrm{III}}=15.1$). This divergence confirms that the class-dependent dynamical signatures emerge specifically from the interaction between polynomial structure and chaotic dynamics; they are not artifacts of the invariant construction itself.
VII.3 Algebraic Mechanism of Class III Spikes
The statistical analysis reveals that Class III invariants exhibit significantly higher intermittency in the chaotic regime. We identify the algebraic mechanism responsible.
The evolution function $Q_{n}(x,y,z)=-\dot{P}_{n}$ is polynomial in $(x,y,z)$; it contains no singular factors. The spikes in $Q_{\mathrm{III}}(t)$ arise not from singularities but from the rapid variation of $\dot{P}_{\mathrm{III}}$ as the trajectory traverses specific regions of phase space.
Two surfaces must be carefully distinguished:
  1. 1.
    Regularization surface (z-nullcline): $p_{\mathrm{III}}=xy-\beta z=0$, where the original auxiliary variable $u_{n}$ diverges.
  2. 2.
    Separatrix: the stable manifold of the saddle point at the origin, locally approximated by $x=0$.
These surfaces are generically distinct. The regularization procedure ensures that all invariants are smooth at the z-nullcline. The spikes arise from the rapid variation of $P_{\mathrm{III}}$ at the separatrix.
The Class III polynomial parts contain terms proportional to $xy$. The evolution function therefore includes:
$\frac{d(xy)}{dt}=\dot{x}y+x\dot{y}=\sigma(y-x)y+x[x(\rho-z)-y].$ (73)
At the separatrix $x=0$: $d(xy)/dt=\sigma y^{2}>0$. The product $xy$ evolves from negative through zero to positive during a left-to-right lobe transition.
Sign of spikes. For $K_{13}$ with $P_{13}=y^{2}/2+xyz-xy\rho+x^{2}\sigma/2-xy\sigma$, the time derivative at $x=0$ is:
$\dot{P}_{13}\big|_{x=0}=y^{2}(-1+\sigma z-\sigma\rho-\sigma^{2}).$ (74)
With $\sigma=10$ and $\rho=28$, this yields $\dot{P}_{13}|_{x=0}=y^{2}(10z-381)$. The sign depends on the altitude $z$ at which the separatrix crossing occurs: for $z>38.1$, $\dot{P}_{13}>0$ and hence $Q_{13}<0$; for $z<38.1$, the signs reverse. In the Lorenz attractor, lobe-switching events occur predominantly in the upper region where $z\approx 40$4545. Taking $z\approx 43$ as representative:
$\dot{P}_{13}\big|_{x=0}\approx y^{2}(430-381)=+49\,y^{2}>0.$ (75)
Therefore $Q_{13}=-\dot{P}_{13}\approx-49\,y^{2}<0$, confirming negative spikes (Fig. 2c). The negative skewness $S_{\mathrm{III}}\approx-2.67$ reflects this predominance of negative excursions at topological transitions.
Why Class I exhibits lower intermittency. The Class I invariant $K_{5}$ also contains an $xy$ term ($xy\rho$), yet exhibits lower kurtosis. The resolution lies in the polynomial structure: $P_{5}=y^{2}/2+\beta z^{2}/2-xy\rho$ contains the term $\beta z^{2}/2$, which $P_{13}$ lacks. The $z^{2}$ term creates continuous baseline fluctuations: since $z$ oscillates between approximately 5 and 45 on the attractor, the $z^{2}$ contribution varies by a factor of approximately 80, generating substantial variance in $Q_{5}$ even during quiescent intervals.
The statistical mechanism: kurtosis scales as $\kappa\propto\mu_{4}/(\sigma_{\mathrm{spike}}^{2}+\sigma_{\mathrm{base}}^{2})^{2}$. Adding baseline variance dilutes the prominence of spikes. Class I has larger total variance from the $z^{2}$-driven baseline, reducing kurtosis.
The Class III polynomial $P_{13}$ couples exclusively to $xy$-dependent terms. All significant variation concentrates at separatrix crossings. Between crossings, $Q_{13}$ remains quiescent. This concentration produces the heavy-tailed distributions characteristic of high kurtosis.
The elevated intermittency of Class III thus reflects not the mere presence of $xy$ terms (which Class I also possesses), but the absence of smoothing terms like $z^{2}$ that distribute variance across the trajectory.
The evolution function
$Q_{\mathrm{III}}(t)=-\dot{P}_{\mathrm{III}}(x(t),y(t),z(t))$ (76)
evaluated along trajectories inherits the rapid variations. The spike magnitude scales as:
$|Q_{\mathrm{III}}|_{\text{max}}\sim\frac{\Delta P_{\mathrm{III}}}{\tau_{\text{cross}}}.$ (77)
Numerical analysis confirms $|Q_{\mathrm{III}}|$ reaches $10^{4}$$10^{5}$ during lobe-switching events.
Distinction from singularities. The regularization procedure ensures that $K_{\mathrm{III}}$ is well-defined throughout phase space. The spikes occur at the separatrix ($x=0$), which is generically not the z-nullcline. The evolution function $Q_{\mathrm{III}}$ is smooth everywhere; the “spikes” are large values of a smooth function, not singularities. The high intermittency reflects geometric sensitivity to topological transitions, not numerical artifacts.
VII.4 Class III Spikes as Symbolic Markers
The Class III evolution function $Q_{\mathrm{III}}(t)$ functions as a symbolic marker that signals each addition to the trajectory’s symbolic sequence.
In the standard symbolic dynamics description [6], each visit to the left lobe is encoded as $L$, each visit to the right lobe as $R$. The sequence $\ldots LLRLLRLR\ldots$ captures the trajectory’s itinerary through the attractor. Each lobe-switching event appends a new symbol to this sequence.
The evolution function $Q_{\mathrm{III}}(t)$ provides a continuous encoding of this discrete dynamics:
  • Quiescent intervals ($|Q_{\mathrm{III}}|\ll|Q_{\mathrm{III}}|_{\text{max}}$): The trajectory remains within a single lobe; no symbols are added.
  • Spike events ($|Q_{\mathrm{III}}|\sim|Q_{\mathrm{III}}|_{\text{max}}$): The trajectory crosses the separatrix; a new symbol is appended.
This establishes $Q_{\mathrm{III}}(t)$ as a geometric probe of topological transitions: the integrated area under each spike corresponds to the “topological contribution” of that symbol transition. Class III invariants thereby encode the discrete symbolic structure of the chaotic flow, while Class I invariants probe the continuous dynamics.
Figure 2: Statistical signatures of the regularization class structure. Probability density functions of standardized evolution functions $Q_{\mathrm{I}}$ (Class I, solid blue) and $Q_{\mathrm{III}}$ (Class III, dashed red) on logarithmic scale. (a) Pre-chaotic regime ($\rho=23$): distributions are statistically indistinguishable ($\kappa\approx 37$ for both). (b) Chaotic regime ($\rho=28$): Class III develops heavier tails ($\kappa_{\mathrm{III}}\approx 15$ versus $\kappa_{\mathrm{I}}\approx 8$), indicating enhanced intermittency. (c) Temporal dynamics in the chaotic regime: $x(t)$ (blue) and $Q_{\mathrm{III}}(t)/5000$ (red, rescaled for visibility). Negative spikes in $Q_{\mathrm{III}}$ correlate with zero-crossings of $x(t)$ (vertical dashed lines marking separatrix crossings). The spikes arise from rapid variation of xyxy-terms at the separatrix; the invariant $K_{\mathrm{III}}$ remains constant throughout.
VIII Discussion
VIII.1 Principal Result: Class-Dependent Dynamical Sensitivity
The central result of this work is that different regularization classes probe different aspects of chaotic dynamics. This distinction transcends formal mathematical structure and reveals operational differences in how the invariants respond to trajectory evolution.
Class I invariants as continuous probes. The polynomial parts $P_{\mathrm{I}}$ contain terms like $z^{2}$ that vary continuously throughout the attractor. The evolution function $Q_{\mathrm{I}}=\dot{v}_{\mathrm{I}}$ exhibits moderate intermittency ($\kappa_{\mathrm{I}}\approx 8$), with variance distributed across quiescent and active intervals alike. These invariants track smooth, continuous aspects of the flow.
Class III invariants as topological probes. The polynomial parts $P_{\mathrm{III}}$ couple directly to the bilinear product $xy$, which changes sign at every lobe transition. The evolution function $Q_{\mathrm{III}}$ exhibits large negative spikes ($\kappa_{\mathrm{III}}\approx 15$) precisely at separatrix crossings, with quiescent intervals between transitions. These invariants function as geometric markers of the symbolic dynamics, encoding the discrete topological structure of the attractor.
This distinction is the primary result of the statistical analysis: the regularization class structure, which emerges from algebraic properties of the orthogonality ansatz, maps onto a meaningful partition of dynamical sensitivity. The elevated kurtosis of Class III is a direct consequence of polynomial structure concentrating significant variation into discrete spike events at lobe switches.
VIII.2 Physical Interpretation of the Conservation Laws
The history-dependent invariants occupy a distinctive position in the taxonomy of conservation laws.
Relation to classical conservation laws. In Hamiltonian mechanics, conservation laws arise from continuous symmetries via Noether’s theorem: energy from time-translation invariance, momentum from spatial translation, angular momentum from rotational invariance. These laws constrain motion to hypersurfaces in phase space and enable reduction of effective dimensionality.
The invariants $K_{n}=P_{n}+v_{n}$ share the formal property of constancy along trajectories but differ in origin and interpretation. They do not arise from a variational principle or symmetry of an action functional. Rather, they emerge from the algebraic structure of the Lorenz flow when augmented with auxiliary variables. Conservation is enforced by construction: $\dot{v}_{n}=-\dot{P}_{n}$ by definition.
This non-Noetherian character connects to the broader framework established by Hojman [5], who demonstrated that conservation laws can be constructed without Lagrangian or Hamiltonian formulations. The present invariants represent a concrete realization of this principle for a dissipative chaotic system.
Non-triviality of the construction. A natural question is whether this conservation is merely tautological. The existence of the null class (Theorem 1) provides a negative answer. If conservation could be achieved for arbitrary polynomials $P(x,y,z)$ simply by defining $v=-\int\dot{P}\,dt$, then all 24 permutations would yield valid invariants. The failure of six permutations demonstrates that Lorenz structure genuinely constrains which polynomial forms are compatible with the orthogonality ansatz.
The null class failure reflects Schwarz integrability conditions imposing algebraic relations among partial derivatives that are violated when the auxiliary variable occupies the leading position. This obstruction is specific to the Lorenz vector field and would differ for other polynomial systems.
Main results. This construction yields two principal results:
First, the classification of compatible polynomial structures encodes information specific to the Lorenz dynamics. Eighteen valid invariants organized into three regularization classes reflect algebraic compatibility between the Lorenz flow and the orthogonality ansatz. Different classes correspond to different nullcline structures, connecting abstract algebraic classification to geometric organization of the flow.
Second, the distinct dynamical signatures exhibited by different classes provide complementary probes of attractor geometry. Statistical analysis demonstrates that Class III invariants function as geometric markers of topological transitions (lobe switches), while Class I invariants track continuous dynamics. This class-dependent sensitivity makes the classification a diagnostic tool for characterizing chaotic trajectories.
Operational interpretation. The invariants are most properly understood as constraints relating instantaneous configuration to accumulated trajectory history. While the form of conservation is guaranteed by $\dot{v}_{n}=-\dot{P}_{n}$, the existence of compatible polynomial structures $P_{n}$ reflects specific algebraic properties of the Lorenz equations.
VIII.3 Regularized Variables: Physical Interpretation
The central insight is that each invariant $K_{n}$ requires its own regularized variable $v_{n}(t)$, even when invariants belong to the same regularization class. Different conservation laws impose different demands on how trajectory history must be accumulated to maintain conservation.
The regularized variable $v_{n}(t)$ can be interpreted as the variation required to conserve $K_{n}$ along a trajectory. The evolution equation $\dot{v}_{n}=Q_{n}(x,y,z)$ specifies how this variation accumulates as the system evolves. The algebraic structure ensures that $v_{n}(t)=c_{n}-P_{n}$ fluctuates within a bounded range determined by attractor geometry. This variation is kinematic in nature, reflecting geometric accumulation of history required to satisfy the constraint $K_{n}=\mathrm{const}$. Unlike thermodynamic quantities (work, entropy production), $v_{n}$ carries no energetic content and does not satisfy a fluctuation-dissipation relation; it functions as a deterministic history accumulator, as discussed in Sec. II.2.
Statistical analysis (Sec. VII) quantifies the fluctuation patterns. Class III invariants incur large-amplitude spikes in $Q_{\mathrm{III}}$ at lobe-switching events, while Class I invariants accumulate more smoothly. This distinction reflects genuine differences in how invariants probe dynamics: Class III invariants are geometrically sensitive to topological transitions, while Class I invariants track more continuous aspects of the flow.
VIII.4 Relation to the Original Invariant
The invariant $K$ presented in Ref. [1] was derived using an orthogonality ansatz of the form $\nabla C=(-\dot{y},-\dot{u},\dot{z},-\dot{x})$, which assigns independent signs $\epsilon_{i}\in\{+1,-1\}$ to each permuted flow component. This represents one element of a larger family: allowing each of the four components to carry an independent sign yields $24\times 2^{4}/2=192$ distinct ansätze (the factor of $1/2$ accounts for the global gauge symmetry $K\to-K$). The systematic exploration presented here restricts attention to the 24 ansätze with uniform global sign, i.e., $\nabla C=\pm\pi(\mathbf{f})$.
Explicit mapping. The ansatz of Ref. [1] corresponds to the permutation $(y,u,z,x)\to(2,4,3,1)$ with sign pattern $(\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4})=(-1,-1,+1,-1)$. However, the regularization polynomial determines class membership, not permutation indices. The original invariant $\mathcal{K}$ has the form (Eq. 8 of Ref. [1]):
$\mathcal{K}=xy(z+1)+\tfrac{1}{2}x^{2}(z-\rho)-\tfrac{1}{2}\beta z^{2}+\sigma v,\quad v=(x-y)u.$ (78)
The regularized variable v=(xy)uv=(x-y)u absorbs the singularity at $x=y$, identifying the regularization polynomial as $(x-y)=-p_{\mathrm{I}}$. Therefore, the original invariant $\mathcal{K}$ belongs to Class I, sharing singularity structure with $K_{1}$$K_{6}$. Despite the non-uniform sign pattern, $\mathcal{K}$ exhibits the same geometric sensitivity as Class I invariants: its evolution function $Q=\dot{v}$ varies smoothly across the attractor without the large spikes characteristic of Class III. This confirms that the regularization class structure, determined by which nullcline removes the singularity, is more fundamental than the specific permutation or sign convention.
Both families share the same regularization structure: the singularity-removing polynomials $p_{\mathrm{I}}=y-x$, $p_{\mathrm{II}}=y+x(z-\rho)$, and $p_{\mathrm{III}}=xy-\beta z$ appear in both constructions. This shared structure suggests that the class organization is a robust feature of the Lorenz system.
VIII.5 Operational Interpretation and Finite Memory Horizons
The history-dependent nature of the invariants raises a practical question: what is the operational meaning of a conserved quantity that depends on the infinite past of a trajectory?
The regularized variable $v_{n}(t)$, defined through $v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}\,ds$, accumulates trajectory history from $t=0$. In the deterministic setting, this accumulation is exact and $K_{n}$ remains conserved. However, in any physical realization, small perturbations (numerical errors, thermal fluctuations, external noise) cause trajectory deviations. Since $v_{n}(t)$ integrates over all time, even small persistent perturbations can lead to secular drift.
A natural resolution interprets the invariants as cycle-averaged quantities. Consider a Poincaré section $\Sigma$ defined by a transverse surface, for instance $z=\rho-1$. Each time the trajectory crosses $\Sigma$, we reset $v_{n}\to 0$. The value accumulated during one return defines production of history for that cycle:
$\Delta v_{n}=\oint_{\text{cycle}}Q_{n}(x,y,z)\,dt.$ (79)
This cycle-averaged interpretation admits a rigorous foundation. For UPOs, $\Delta v_{n}$ takes a fixed value characteristic of the orbit, providing the same fingerprint discussed earlier. For chaotic trajectories, $\{\Delta v_{n}\}$ forms a discrete-time observable whose statistical properties are governed by the ergodic structure of the attractor.
VIII.6 Robustness Under Stochastic Perturbations
The deterministic conservation laws $dK_{n}/dt=0$ hold exactly when the system evolves according to the Lorenz equations (1)–(3). In the presence of additive noise, a more realistic model of physical systems, the evolution becomes
$dx$ $=\sigma(y-x)\,dt+\epsilon\,dW_{x},$ (80)
$dy$ $=[x(\rho-z)-y]\,dt+\epsilon\,dW_{y},$ (81)
$dz$ $=(xy-\beta z)\,dt+\epsilon\,dW_{z},$ (82)
where $dW_{i}$ are independent Wiener increments and $\epsilon$ measures noise strength.
Under such perturbations, the exact cancellation that ensures $dK_{n}/dt=0$ fails. The invariant $K_{n}$ becomes a stochastic process:
$dK_{n}=\epsilon\sum_{i}\frac{\partial K_{n}}{\partial x_{i}}dW_{i},$ (83)
a random walk with variance growing as $\epsilon^{2}t$.
The characteristic time over which invariants remain approximately conserved scales as:
$\tau_{\text{useful}}\sim\frac{1}{\epsilon^{2}\langle\|\nabla_{xyz}K_{n}\|^{2}\rangle_{\mathrm{SRB}}},$ (84)
where $\langle\cdot\rangle_{\mathrm{SRB}}$ denotes the average over the SRB measure. The quadratic sensitivity to noise distinguishes history-dependent invariants from topological invariants (such as knot types of periodic orbits), which are robust under arbitrary continuous deformations.
Diagnostic utility of structural redundancy. The existence of 18 algebraically distinct invariants with functional rank 4 provides internal consistency checks. For any pair of invariants within the same class, the identity $v_{n}(t)-v_{m}(t)=P_{m}(x,y,z)-P_{n}(x,y,z)$ must hold exactly in the deterministic limit, and to within Brownian diffusion under stochastic perturbations. A sudden violation of this identity, beyond the expected $O(\epsilon\sqrt{t})$ drift, signals numerical integration failure rather than physical dynamics. This is particularly valuable near nullcline crossings where stiffness is high. Furthermore, comparing invariants from different classes provides complementary information: Class III invariants signal topological transitions (lobe switches), while Class I invariants track continuous dynamics. Under noise, a spike in $Q_{\mathrm{III}}$ accompanied by smooth $Q_{\mathrm{I}}$ behavior confirms a genuine separatrix crossing, whereas uncorrelated fluctuations in both suggest stochastic perturbation. The structural redundancy thus functions as a filter bank for distinguishing topology from noise.
VIII.7 Numerical Stiffness and Practical Integration
While the regularization procedure (Sec. III.4) ensures that all invariants are theoretically well-defined throughout phase space, practical numerical integration of Class III evolution functions presents challenges.
The evolution equation $\dot{v}_{n}=Q_{n}(x,y,z)$ requires evaluating the polynomial $Q_{n}$ along the trajectory. For Class III invariants, statistical analysis reveals that $|Q_{\mathrm{III}}|$ reaches values of order $10^{4}$$10^{5}$ during lobe-switching events, while remaining of order $10^{0}$$10^{1}$ between transitions. This disparity constitutes numerical stiffness.
The stiffness ratio can be estimated as:
$\mathcal{S}\equiv\frac{|Q_{n}|_{\text{max}}}{|Q_{n}|_{\text{typical}}}\sim\frac{T}{\tau_{\text{cross}}}\approx 10^{3}\text{--}10^{4},$ (85)
where $T\approx 1.5$ is the typical orbital period and $\tau_{\text{cross}}\approx 10^{-3}$$10^{-4}$ is the separatrix crossing time.
Three practical implications follow:
  1. 1.
    Adaptive step control: The StiffnessSwitching method employed in validation automatically detects stiff intervals and adjusts the integration algorithm. This is essential for maintaining accuracy during Class III spikes.
  2. 2.
    Class-dependent accuracy: Class I and Class II invariants, lacking the large spikes of Class III, can be integrated with coarser timesteps. For applications requiring only Class I fingerprints, computational costs are substantially reduced.
  3. 3.
    Spike-aware sampling: When computing statistics of $Q_{n}(t)$, fixed-interval sampling may undersample spike events. The heavy tails in Fig. 2(b) are accurately captured only with sufficiently fine temporal resolution ($\Delta t\ll\tau_{\text{cross}}$).
The distinction between analytical smoothness and numerical stiffness is important: invariants are mathematically well-defined everywhere, but practical computation requires integration schemes capable of handling large derivative ratios. This is a standard challenge in stiff ODE systems and does not reflect any deficiency in the invariant construction.
VIII.8 Forward and Inverse Problems
A complete assessment of the invariants’ utility requires distinguishing two complementary perspectives.
The Forward Problem. In the forward or initial value problem, one specifies initial conditions $(x_{0},y_{0},z_{0})$ and asks: where does the trajectory go? Since $v_{n}(0)=0$ for all $n$, the invariant values are determined instantaneously:
$c_{n}=K_{n}(0)=P_{n}(x_{0},y_{0},z_{0}).$ (86)
The Triad fingerprint $(c_{1},c_{2},c_{3})$ is determined algebraically from initial conditions without integration. The fingerprint is an a priori property of the trajectory.
The Inverse Problem. In the inverse or classification problem, one specifies a Triad of invariant values $(c_{1},c_{2},c_{3})$ and asks: what initial conditions are compatible? The three polynomial equations $P_{n}(x,y,z)=c_{n}$ generically determine a discrete set of candidate initial states. This provides a classification tool: periodic orbits can be catalogued by fingerprints, and chaotic trajectories can be organized into families based on invariant values.
VIII.9 From Algebra to Dynamics
The partition into three classes based on regularization polynomial $p_{n}$ was established algebraically. Statistical analysis demonstrates genuine dynamical consequences, elevating the classification from mathematical convenience to physically meaningful distinction.
The class structure reflects different response of $Q_{n}(t)$ to trajectory motion:
  • Class I (regularized by $p_{\mathrm{I}}=y-x$): The nullcline $y=x$ is crossed at various orbital phases, not specifically at lobe-switching events. This explains moderate kurtosis ($\kappa_{\mathrm{I}}\approx 8$).
  • Class III (regularized by $p_{\mathrm{III}}=xy-\beta z$): Polynomial parts contain terms proportional to $xy$, which changes sign at separatrix crossings. Class III invariants are sensitive to topological transitions. Elevated kurtosis ($\kappa_{\mathrm{III}}\approx 15$) and strong negative skewness ($S_{\mathrm{III}}\approx-2.7$) reflect the concentration of large negative spikes at lobe switches, where transitions occur in the upper region of the attractor.
The spikes in $Q_{\mathrm{III}}(t)$ function as markers signaling each addition to the symbolic sequence $\ldots LLRLLRLR\ldots$. Class III invariants serve as geometric probes of topological transitions.
IX Conclusion
This work has presented a systematic exploration of history-dependent dynamical invariants in the Lorenz system using permutation-based enumeration. The principal results are:
1. Enumeration: Of 24 possible permutation ansätze with uniform global sign, exactly 18 yield valid invariants ($K_{1}$$K_{18}$). The failure of 6 permutations demonstrates that Lorenz structure constrains which polynomial forms can be conserved; the construction is algebraically constrained, not arbitrary.
2. Classification: The invariants organize into three classes (I, II, III) determined by the regularization polynomial, plus a null class (4abc) whose failure reflects Schwarz integrability conditions of the augmented system.
3. Regularized variables: Each invariant $K_{n}$ requires its own regularized variable $v_{n}(t)$ satisfying $\dot{v}_{n}=Q_{n}(x,y,z)$. Regularized variables remain bounded, as established by $v_{n}=c_{n}-P_{n}$ and compactness of the attractor. Original auxiliary variables $u_{n}=v_{n}/p_{n}$ diverge at nullcline crossings, representing coordinate singularities that do not affect invariant conservation.
4. Independence of history constraints: Any selection of three invariants from different regularization classes provides independent constraints on trajectory histories. The Triad fingerprint $(c_{1},c_{2},c_{3})$ characterizes each trajectory up to the discrete $\mathbb{Z}_{2}$ symmetry. Parity-breaking Triads (containing odd invariants) can distinguish symmetric trajectories.
5. Immediate determination: The Triad fingerprint is determined instantaneously from initial conditions via $c_{n}=P_{n}(x_{0},y_{0},z_{0})$, since $v_{n}(0)=0$. The constraints are a priori properties of trajectories.
6. Class dynamics: The three regularization classes exhibit distinct dynamical signatures in the chaotic regime. Class III displays higher intermittency (kurtosis $\kappa_{\mathrm{III}}\approx 15$ versus $\kappa_{\mathrm{I}}\approx 8$; Fig. 2b) and stronger negative skewness ($S_{\mathrm{III}}\approx-2.7$ versus $S_{\mathrm{I}}\approx-1.6$), correlating with lobe-switching events in the upper attractor region. This divergence is absent pre-chaos (Fig. 2a), confirming it arises from the interaction between polynomial structure and chaotic dynamics. Class III evolution functions $Q_{\mathrm{III}}(t)$ exhibit negative spikes (Fig. 2c) that encode the symbolic dynamics of the chaotic flow.
7. Physical interpretation: The regularized variable $v_{n}(t)$ represents the kinematic variation required to conserve $K_{n}$ along a trajectory. Different invariants have different variation functions, and the statistical properties of these variations distinguish continuous-tracking (Class I) from topological-transition-sensitive (Class III) invariants.
8. Utility: The invariants provide: (i) independent monitors of numerical integration accuracy, (ii) unique fingerprints for classifying unstable periodic orbits, (iii) trajectory-discriminating quantities that encode information about the initial state not recoverable from the current position alone, (iv) complementary probes for characterizing attractor structure, with Class III invariants serving as geometric probes of topological transitions while Class I invariants track continuous dynamics, and (v) a conceptual framework relating chaos to incompatibility of different measures of trajectory history.
9. Relation to original invariant: The invariant $\mathcal{K}$ of Ref. [1] employs a non-uniform sign ansatz belonging to the complementary 192-element family. The regularized variable v=(xy)uv=(x-y)u identifies the regularization polynomial as $(x-y)=-p_{\mathrm{I}}$, placing $\mathcal{K}$ in Class I. This confirms that class structure is determined by the nullcline that removes the singularity, not by permutation indices or sign pattern.
Open questions: Several problems emerge from this analysis: (i) What is the complete structure of the 192-element family with independent component signs? (ii) Does this framework extend to other dissipative chaotic systems, and what determines invariant family structure? (iii) Can the bounded fluctuations of regularized variable differences $v_{n}(t)-v_{m}(t)$ be exploited as diagnostics of attractor geometry? (iv) What is the explicit form of the probability density function of cycle-averaged $\Delta v_{n}$, and how does its structure relate quantitatively to attractor symbolic dynamics? (v) In stochastic settings, can dynamic selection among invariants with locally minimal $\|\nabla P_{n}\|$ extend the useful prediction horizon beyond the estimate (84)?
The systematic framework established here, comprising permutation-based enumeration with regularization analysis and statistical characterization, is applicable to other dissipative systems and may reveal similar hidden structure in chaotic dynamics.
Acknowledgments
This research was conducted independently at the Departamento de Física, Universidad de Chile, without external funding.